tiaunabuchan:

Use the function f(x) = 2x3 - 3x2 + 7 to complete the exercises. help

XioGonz:

f(−1) =2 f(1) =6 f(2) =11

tiaunabuchan:

thank u

CocoCammron:

f(−1) = ✔ 2 f(1) = ✔ 6 f(2) = ✔ 11

justus:

It is against QC rules to give direct answers without guidance of the user so here is an explanation on how they supposedly got the answer(s). \[1. f(-1)=2*(-1)*3-3*(-1)^2+7=-2-3+7=2\] \[2. f(1)=2*1^3-3*1^2+7=2-3+7=6\] \[3. f(2)=2*2^3-3*2^2+7=16-12+7=11\]